OpenStudy (anonymous):

solve the following quadratic equation by using the quadratic formula 3x^2+7x+2=0

OpenStudy (jiteshmeghwal9):

given: a=3 b=7 c=2 \[-(7) \pm \sqrt{(7)^2-4(3)(2)}\over2(3)\]

OpenStudy (jiteshmeghwal9):

\[-7 \pm \sqrt{49-24}\over6\]

OpenStudy (jiteshmeghwal9):

\[-7 \pm \sqrt{25}\over 6\]

OpenStudy (jiteshmeghwal9):

\[-6 \pm 5 \over6\]so\[\alpha={-6+5\over6},\beta={-6-5\over6}\]solve this two roots & gt ur answer:)

OpenStudy (anonymous):

Thank you so much thank you for your response, i understand it a lot better! :)